Please help me solve this equation:
Subtract and simplify
7/9-11/18=

Found 2 solutions by nadinadan, i_dontget_prealgebra:
Answer by nadinadan(32)

7/9-11/18 So we need to have the same denominators on both fractions.
The common denominator is 18, so we have to multiply the first fraction by 2.
so 14/18-11/18=(14-11)/18=3/18 simplified we can divide the nominator and denominator by 3 so the answer is 1/6.
Let me know if you need more help.

Answer by i_dontget_prealgebra(1)

7/9-11/18 So we need to have the same denominators on both fractions.
The common denominator is 18, so we have to multiply the first fraction by 2.
so 14/18-11/18=(14-11)/18=3/18 simplified we can divide the nominator and denominator by 3 so the answer is 1/6.
Glad i could help. im surprised i even answered this one. i usually dont understand these